OpenStudy (anonymous):

A rectangular board is measured to have a width, w, of 1.25 ± 0.01 m and length, l, of 9.58 ± 0.02 m . What is the absolute uncertainty of the width? Also What is the perimeter, p, of the above board and the the experimental uncertainty in the perimeter, δ p?

OpenStudy (anonymous):

P = 2 (L+B) = 2( 1.25 + 9.58) uncertainty = P +- (0.01+0.02)